Complete in Excel show the formula please if applicable

Three options are being considered by the human resources management of Suncore Distribution Center to manage a new distribution warehouse. The following payoff table gives the profits ($000) possible for the different options and possible future demand levels.

State of Nature (Demand level)

Alternatives

High           

Medium          

Low                  

Use own staff

650

550

625

Outside vendor

900

600

400

Combination

800

750

500

a. What would the recommendation be if the optimistic (maximax) point of view was taken?

b. What would the recommendation be if the pessimistic (maximin) point of view was taken?

c. What would the recommendation be if the Equal Likelihood method was used?

d. Assume the following probabilities are estimated for the future market share

                        high .25                        medium .45                 low .30

            Calculate the Expected monetary values for the three options and give the recommendation based on EMV.

e. Calculate the Expected value with Perfect Information and the Expected value of Perfect Information             (assuming probabilities given in part d).

f. Interpret the EV of PI
Answer #1

a) Maximax means take Maximum of Maximum of all alternatives i.e Max of (650, 900,800) i.e. 900. So outside vendor will be chosen.

b) Maximin means take best of the worst output i.e. Max (550,400,500) . With this we will decide to use own staff.

c) Equal likelihood means each state has equal probability.

Payoff on using own staff = (650+550+525) / 3 =575

Payoff on using outside vendor = (900+600+400) / 3 =633.33

Payoff on using oCombination = (800+750+500) / 3 = 683.33

We will choose combination model.

d) Expected payoff = Sum of (Payoff * Probability of state)

For Combination = 800*0.25 + 750*0.45+500*0.30 = 687.50 (Maximum)

Alternatives High            Medium           Low                  
Use own staff 650 550 625
Outside vendor 900 600 400
Combination 800 750 500
Probability 0.25 0.45 0.3 Total
Use own staff 162.50 247.50 187.50 597.50
Outside vendor 225.00 270.00 120.00 615.00
Combination 200.00 337.50 150.00 687.50

e) Expected value without perfect information is = 687.50 (The decision we will take given probabilities)

The expected value with Perfect Information = Output when you know the probabilities possible with surety = 900*0.25+750*0.45+625*0.30 = 750

Expected value of perfect information = expected value with Perfect Information-Expected value without perfect information = 750 - 687.5 = $62.5

f) If we can know the probability of each state (perfect information) in $62.5 or less we should consider it. However, if the value of information is more than $62.5 then we must not consider taking it.